OpenStudy (anonymous):

Find the equation of the tangent line y= 1/cosx - 2cosx at (pi/3, 1)

OpenStudy (anonymous):

I think i did my derivative wrong i got 2 tan^2 (x) / cosx

OpenStudy (zzr0ck3r):

take the derivitave then plug in your x value to get slope, then use y = mx+b with your one point to figure out b

OpenStudy (anonymous):

I know what to do , but i just got the answer wrong..

OpenStudy (anonymous):

what's your slope?

OpenStudy (anonymous):

Shouldn't you have two different expressions?

OpenStudy (anonymous):

the slope is just the derivative 2 tan^2 (x) / cosx umm..i always just use one..what u mean by 2 different expressions?

OpenStudy (anonymous):

cuz i took the common denominator

OpenStudy (anonymous):

evaluate your derivative at x=pi/3

OpenStudy (anonymous):

I did, but i got a completely different answer than the book's answer so i am thinking i might do the derivative wrong can someone show me how to get the derivative

OpenStudy (zzr0ck3r):

dy/dt = (sin(x)/cos^2(x))(2cos(x)^2+1)

OpenStudy (anonymous):

Just wolfram it if you are checking your work http://www.wolframalpha.com/input/?i=+1%2Fcosx+-+2cosx+derivative

OpenStudy (anonymous):

what did you get when you put pi/3 into your derivative?

OpenStudy (anonymous):

how u get (2cos(x)^2+1), and shouldnt it be a + between

myininaya (myininaya):

\[y=\frac{1}{\cos(x)} -2\cos(x) ?\]

OpenStudy (zzr0ck3r):

d/dt(1/cosx) = sin(x)/cos(x)^2 this is from the quotent rule d/dt(-2cos(x) ) = 2sin(x)

OpenStudy (anonymous):

is that your function? (the one @myininaya typed?) that's what i used.

OpenStudy (anonymous):

yes

OpenStudy (zzr0ck3r):

then add them together

myininaya (myininaya):

\[y=\sec(x)-2\cos(x) => y'=\sec(x)\tan(x)+2\sin(x)\]

OpenStudy (zzr0ck3r):

the answer can look 1000 different ways ....

OpenStudy (anonymous):

yea, i did what myininaya shows , but i further took the common denominator ...so i guess i should not do it?

OpenStudy (anonymous):

but there is only one slope at x=pi/3

OpenStudy (zzr0ck3r):

no dont take the common den it will make it harder

OpenStudy (zzr0ck3r):

just split it into two derivatives

OpenStudy (zzr0ck3r):

d/dt(x+y) = dx/dt+dy/dt

OpenStudy (anonymous):

ohh ..i will try to do the question again without spilting them together thankss

myininaya (myininaya):

\[y'_{(\frac{\pi}{3},1)}=\sec(\frac{\pi}{3}) \tan(\frac{\pi}{3}) +2 \sin(\frac{\pi}{3})\] So\[\cos(\frac{\pi}{3})=\frac{1}{2} ; \sin(\frac{\pi}{3})=\frac{\sqrt{3}}{2}\]

OpenStudy (zzr0ck3r):

I think his problem is calculating the derivative

myininaya (myininaya):

oh i thought he said he got what i got?

OpenStudy (anonymous):

it is calculate the equation of the tangent

OpenStudy (anonymous):

so the derivative is the slope and solve for b and then y and x..make it in standard form like that

OpenStudy (anonymous):

you'll have to get the derivative first. and whether it is simplified or not, plugging in x=pi/3 into will give you the slope.

OpenStudy (anonymous):

because that's all you need. you have a point, you have slope... put it in point slope form.

OpenStudy (zzr0ck3r):

he knows all this, the question should of been "how do i take this derevative" i think

OpenStudy (anonymous):

ok thanks everyone, i think i will know how to figure out the rest on my own openstudy sucks that can only give out 1 medal :/

OpenStudy (anonymous):

:)

OpenStudy (zzr0ck3r):

lol

myininaya (myininaya):

imaginary medals for everyone lol

OpenStudy (anonymous):

thanks everyone!!!!!! so much so much !!! really dont know who to give lol imaginery medals everyone! thanks